# Local Advice and Local Decompression

### v1

@@ 4. LCLs on Graphs with Subexponential Growth
In this section, we show that on graphs of subexponential growth, we can solve any LCL problem in constant time by using $1$ bit per node of advice (we remind the reader that, as per definition of LCLs, the class of graphs that we consider are the ones with constant maximum degree, hence throughout this section we assume $\Delta=O(1)$). More in detail, we will prove the following theorem, and devote the rest of the section to proving it.
Theorem 1.
@@
Let $\Pi$ be an LCL problem, and let $\mathcal{G}$ be a family of graphs of maximum degree $\Delta$ of subexponential growth in which $\Pi$ is solvable. Then, there exists a uniform fixed-length sparse $(\mathcal{G},\Pi,1,O(1))$-advice schema.
@@
We start by giving a formal definition of subexponential growth. For a node $v$, we denote with $N_{\leq d}(v)$ the set of nodes at distance at most $d$ from $v$, and we denote with $N_{=d}(v)$ the set of nodes at distance exactly $d$ from $v$.
Definition 0 (Subexponential growth).
@@
A family $\mathcal{G}$ of graphs is of subexponential growth if, for any constant $c>0$, there exists a constant $x_{0}$ such that, for any $x\geq x_{0}$, for any $G=(V,E)\in\mathcal{G}$, for any $v\in V$, $|N_{\leq x}(v)|\leq 2^{c\cdot x}$.
@@
On a high level, we would like to exploit that, on graphs with subexponential growth, it is possible to somehow compute a clustering such that the nodes in the border of each cluster are much less than the nodes inside each cluster. Ideally, this property would allow us to encode the solution of the nodes of the border of a cluster by using the nodes inside that cluster. Then, inside each cluster, we can complete the solution by brute force, and hence each node would know its own part of the solution, as desired. However, things are not so simple: (1) not only we need to encode the solution of the nodes that are in the borders of the clusters, but we also need to be able to encode a clustering of the graph, and all this by using only one bit per node; (2) we need to make sure that the decoding process is able to decode the right thing, and this requires that the nodes that encode some information and that are inside two different clusters are far-enough from each other. This latter issue forces us to prove some stronger properties of graphs with subexponential growth, that is that it is possible to compute a clustering such that the number of nodes that are very close to the center of the cluster is at least as much as the number of nodes in the border (Lemma 3). More in detail, we prove that, for each node $v$, there exists some value $\alpha$ such that, for a cluster $C_{v}$ centered at $v$, the set $N_{\leq\alpha}$ of nodes of $C_{v}$ that are at distance at most $\alpha$ from $v$ is such that all nodes in $N_{\leq\alpha}$ are far-enough from the border and they are at least as many as the ones in the border. We will use only nodes in $N_{\leq\alpha}$ to encode any useful information for us. In order to encode the clustering we do the following. First, we compute a distance-$d$ coloring of the graph (for some fixed $x$), that is, a coloring of the nodes such that nodes with the same color are at distance strictly larger than $d$. Then, we proceed through color classes in some order and at each phase $i$ we assign nodes to some cluster centered at a node with color $i$ (the nodes that will not be assigned to any cluster in this phase will be processed later). In this case, we say that $i$ is the color of the cluster. After we are done with color-class $i$, we remove the nodes that are part of some cluster and we repeat the procedure by considering the next color on the remaining subgraph. The distance coloring ensures that we avoid any kind of collisions with other centers of the same color. In order to encode the color of a cluster $C_{v}$ centered at $v$ we use a path that starts at node $v$ and is contained in $N_{\leq\alpha}$ (and hence it is far-enough from the border of the cluster). Then, we use an independent set of the nodes in $N_{\leq\alpha}$ in order to encode the solution of the border (Lemma 3 ensures that these nodes are enough for our purposes). The fact that we use only nodes in $N_{\leq\alpha}$ to encode information ensures that nodes that carry information regarding different clusters are far-enough from each other (even if the clusters are adjacent) and hence we avoid wrong decodings. In order to be able to distinguish different encodings inside a cluster (e.g., the color of the cluster, the output of the border, the center of the cluster) we use the following scheme. Consider the connected components induced by the nodes having as input a $1$: the $1$s that form an independent set are the ones that encode the solution of the border; the other $1$s are used to encode the cluster center and its color, where four $1$s encode the center of the cluster, two $1$s encode a 00 and three $1$s encode a $1$.
@@
We start by proving a technical lemma regarding graphs of subexponential growth, which will be useful later.
Lemma 0.
@@
Let $\mathcal{G}$ be a family of graphs of subexponential growth that have degree bounded by $\Delta$. For any integer $r>0$, let $c:=\log(1+1/\Delta^{r})/(3r)$, and let $x:=\max\{4r,x_{0}\}$, where $x_{0}$ is the value given by Definition 2 for $\mathcal{G}$, when using parameter $c$. Then, for all $G=(V,E)\in\mathcal{G}$, and all $v\in V$, there exists a value $\alpha$ in $\{x,\ldots,2x\}$ such that $|N_{\leq\alpha}(v)|\geq\Delta^{r}|N_{=\alpha+r}(v)|$.
Proof.
@@
Assume, for a contradiction, that, for any $\alpha$ in $\{x,\ldots,2x\}$, $|N_{\leq\alpha}(v)|<\Delta^{r}|N_{=\alpha+r}(v)|$. We prove, by induction, that, under this assumption, $|N_{\leq x+kr}(v)|\geq(1+1/\Delta^{r})^{k}\cdot|N_{\leq x}(v)|$, for any $k$ satisfying $0\leq kr\leq x$. For the base case $k=0$, we need to prove that $|N_{\leq x}(v)|\geq|N_{\leq x}(v)|$, which trivially holds. Assume that the statement holds for $k$, we show that it holds for $k+1$. We can lower bound $|N_{\leq x+(k+1)r}(v)|$ with $|N_{\leq x+kr}(v)|+|N_{=x+(k+1)r}(v)|\geq|N_{\leq x+kr}(v)|+|N_{\leq x+kr}(v)|/\Delta^{r}=|N_{\leq x+kr}(v)|(1+1/\Delta^{r})\geq(1+1/\Delta^{r})^{k}\cdot|N_{\leq x}(v)|\cdot(1+1/\Delta^{r})=(1+1/\Delta^{r})^{k+1}\cdot|N_{\leq x}(v)|$, proving the inductive step. Moreover, by the assumption of subexponential growth, we have that $|N_{\leq x+kr}(v)|\leq 2^{c(x+kr)}$. By combining the obtained bounds, we obtain the following.
$(1+1/\Delta^{r})^{k}\cdot|N_{\leq x}(v)|\leq|N_{\leq x+kr}(v)|\leq 2^{c(x+kr)}.$
@@
$(1+1/\Delta^{r})^{k}\cdot|N_{\leq x}(v)|>2^{c(x+kr)}.$
Since $|N_{\leq x}(v)|\geq 1$, it is sufficient to show that there exists a value of $k$ satisfying $0\leq kr\leq x$, such that
$(1+1/\Delta^{r})^{k}>2^{c(x+kr)},\text{ which is implies by }k>\frac{x}{\log(1+1/\Delta^{r})/c-r}.$
Since $c=\log(1+1/\Delta^{r})/(3r)$, we get that, in order to get our contradiction, it is sufficient to give an integer $k$ satisfying $k>x/(2r)$ and $0\leq kr\leq x$. For this purpose, we take $k:=\lceil\frac{x}{2r}\rceil+1$. Observe that $k>x/(2r)$ is clearly satisfied. Also, $kr\leq(\lceil\frac{x}{2r}\rceil+1)r\leq(\frac{x}{2r}+2)r=x/2+2r\leq x$, where the last inequality holds since $x\geq 4r$ (recall that $x$ is defined as $\max\{4r,x_{0}\}$). ∎
@@
In order to show the desired advice schema, we first compute a clustering of the graph as follows. Let $r$ be a large-enough constant to be fixed later. Let $c$ and $x$ be the values given by Lemma 3 as a function of $r$. Let $G\in\mathcal{G}$. At first, we compute a distance-$(5x)$ coloring on $G$. Since $G$ is a graph of subexponential growth, such a distance coloring can be done by using $2^{c\cdot 5x}$ colors (this is because $x\geq x_{0}$, where $x_{0}$ is the value given by Definition 2 when using parameter $c$). Then, we process nodes by color classes, in ascending order, and, at each step $i$, we assign some nodes to some clusters. Let $G_{i}$ be the subgraph of $G$ induced by nodes that, at the end of step $i-1$, do not belong to any cluster yet. Let $v$ be a node of color $i$. We consider two possible cases: either, in $G_{i}$ it holds that $|N_{=2x}(v)|=0$, or in $G_{i}$ it holds that $|N_{=2x}(v)|>0$. For each node in the latter case, we compute a cluster as follows. Let $\alpha_{v}$ be the value of $\alpha$ given by Lemma 3 for node $v$. We create a cluster containing $v$ and all the nodes of $G_{i}$, that, in $G_{i}$, have distance at most $\alpha_{v}+r$ from $v$. We call $v$ the center of this cluster. Observe that, at the end of this cluster-formation procedure, each node $v$ that does not belong to any cluster satisfies that, within distance $2x$, it sees the whole connected component containing $v$ induced by nodes that do not belong to any cluster.
@@
We provide an assignment of bits to the nodes of $G$ that encodes the clustering. We will later show how to add other bits on top of such an assignment and encode a solution of $\Pi$ in such a way that it is clear which bits encode the clustering and which bits encode the solution of $\Pi$. For each cluster $C$, let $v$ be its center, and let $i$ be the color of $v$. For any $u\in C$, let $d_{u}$ be the distance of $u$ from $v$. By the construction of $C$, there must exist $u\in C$ such that $d_{u}=y:=\lfloor x/2\rfloor$. Consider an arbitrary path $P_{v}=(v_{1},\ldots,v_{y})$ of length exactly $y$ where $v_{1}=v$ and $v_{y}=u$. This is the path that we will use for encoding the color of $v$. Let $B$ be the bit-string representation of $i$, which is of length $\lceil\log i\rceil$. We modify $B$ by replacing each 00 with the bit-string $110$, and each $1$ with the bit-string $1110$, obtaining the bit-string $B^{\prime}$. Then, we define $B^{\prime\prime}$ as the concatenation of $11110110$ with $B^{\prime}$ and with 00. In total, we use at most $94|B|+9=4\lceil\log i\rceil+9\leq 4\lceil\log(2^{c\cdot 5x})\rceil+9=20cx+$ 9 bits. By setting $r$ large enough, we obtain a small-enough $c$ and a large-enough $x$ such that $20cx+9\leq\lfloor x/2\rfloor=y=|P_{v}|$. Hence, we can assign the $j$th bit of $B^{\prime\prime}$ to the node $v_{j}$ of $P_{v}$. Finally, we assign the bit 00 to all nodes that did not get any bit from the previous bit-assignment. Note that all nodes within distance $\alpha_{v}+r$ (in $G_{i}$) from $v$ that are not part of $P_{v}$ have their bit set to 00.
@@
Proof.
If $u$ and $w$ have the same color, then they are at distance at least $5x$, and hence all nodes in $N_{\leq\alpha_{u}}(u)$ are at distance at least $x$ from nodes in $N_{\leq\alpha_{w}}(w)$, and hence the property holds. Therefore, assume, w.l.o.g., that $u$ has color $i$ and $w$ has color $j$, for $i<j$. Then, in $G_{i}$, all nodes in $C_{w}\cap N_{\leq\alpha_{w}}(w)$ are at distance at least $r>0$ from all nodes in $C_{u}\cap N_{\leq\alpha_{u}}(u)$, since the cluster $C_{u}$ contains all nodes within distance $\alpha_{u}+r$ from $u$ in $G_{i}$. This implies that they are at distance strictly greater than zero also in $G$. ∎
Corollary 0.
@@
Proof.
The proof follows from Lemma 4 and by the fact that each path contains only nodes within distance $y<\alpha$ from the center of its cluster. ∎
@@
Assume that we already decoded the clusters having centers of color $11,\ldots,i-$ 1. We now show how to decode clusters having centers of color $i$. Each node gathers its $2x$-radius neighborhood. In this way, each node $v$ knows which of the nodes within distance $2x$ from $v$ are part of some cluster with center of color at most $i-1$, and which nodes within distance $2x$ do not belong to any cluster yet. Hence, each node $v$ knows which nodes within distance $2x$ are part of $G_{i}$. In the following, all computation will be done with regard to $G_{i}$.
@@
  • for any $y+1\leq j\leq x$, nodes in $N_{=j}(v)$ have their bit set to 00;
  • there exists a path $P=(v_{1},\ldots,v_{y})$ satisfying that, for all $j$, the node $v_{j}$ is at distance exactly $j-1$ from $v$, and that contains all nodes in $N_{\leq y}(v)$ that have their bits set to $1$;
@@
[…]
By the encoding of the clusters, it is clear that all centers of color $i$ must satisfy these properties. Consider the set $S^{\prime}\subseteq S$ of nodes satisfying that, by decoding the bit-string on the path, they obtain color $i$. We prove that nodes in $S^{\prime}$ are all and only the centers of clusters of color $i$. Recall that, when we encoded clusters having a center $v$ of color $i$, we assigned a 00 to all nodes within distance $\alpha_{v}+r$ (in $G_{i}$) from $v$ that are not part of $P_{v}$. Hence, it is clear that, if a center $v$ has color $i$, then $v\in S^{\prime}$. Let $v$ be an arbitrary node in $S^{\prime}$, and let $P=(v_{1},\ldots,v_{y})$ be a path satisfying the properties above. Then we show that $v$ must be a center of color $i$. By Corollary 5, nodes belonging to different paths form different connected components, and hence, if four nodes form a connected component containing four $1$s, then one of these nodes must be the center of a cluster. This implies that one node among $v_{1},\ldots,v_{4}$ must be the center of a cluster. By the above properties, we get that no node among $v_{2},v_{3},v_{4}$ can be in $S^{\prime}$, and hence $v_{1}=v$ is indeed a center of a cluster. Also, by the above properties, there are no other centers within distance $x$ from $v$. We therefore get that the bit-string encoded on $P$ is exactly the color of $v$, and hence that the color of $v$ is $i$.
@@
Let $\ell$ be a function mapping each node-edge pair of $G$ into a label of $\Pi$, such that the labeling obtained by $\ell$ is a solution for $\Pi$. We process clusters by colors in ascending order. Let $C_{v}$ be a cluster centered at $v$ with color $i$. In the following, all computation is done w.r.t. $G_{i}$. Let $\bar{r}$ be the checkability radius of $\Pi$. Let $S_{v}$ be the set of nodes within distance $\bar{r}$ (in $G_{i}$) from at least one node in $N_{=\alpha_{v}+r}(v)$. Note that $|S_{v}|\leq N_{=\alpha_{v}+r}(v)\Delta^{\bar{r}}$. Hence, by Lemma 3, $|S_{v}|\leq|N_{\leq\alpha_{v}}(v)|\Delta^{\bar{r}}/\Delta^{r}$. Consider the labeling given by $\ell$ to all node-edge pairs incident to the nodes in $S_{v}$. We can encode such a labeling by using a bit-string of length $k|N_{\leq\alpha_{v}}(v)|\Delta^{\bar{r}+1}/\Delta^{r}$, where $k$ is a constant that depends on the alphabet of $\Pi$. Let $B$ be such a bit-string. Let $Y$ be the set containing the nodes of $N_{\leq\alpha_{v}}(v)$ that have their bit set to $1$, plus all their neighbors. Let $Z=N_{\leq\alpha_{v}}(v)\setminus Y$. The nodes that have their bits set to $1$ are at most $y=\lfloor x/2\rfloor$, and each such node has at most $\Delta-1$ neighbors that have a 00. Recall that $v$ has at least one node at distance exactly $2x$ (as otherwise $v$ would not be a cluster center), and that $N_{\leq\alpha_{v}}(v)$ contains all nodes within distance $\alpha_{v}\geq x$ from $v$. We get that there exists an injective function mapping nodes that have a $1$ to nodes in $Z$. Thus, $|Z|\geq|Y|/\Delta$, and hence $|N_{\leq\alpha_{v}}(v)|=|Z|+|Y|\geq|Y|/\Delta+|Y|=|Y|(1+1/\Delta)$, which implies $|Y|\leq|N_{\leq\alpha_{v}}(v)|/(1+1/\Delta)$. Thus, we get that $|Z|\geq|N_{\leq\alpha_{v}}(v)|-|Y|\geq|N_{\leq\alpha_{v}}(v)|/(\Delta+1)$. Hence, any independent set in $Z$ has size at least $|N_{\leq\alpha_{v}}(v)|/(\Delta+1)^{2}$. By picking $r$ large enough, we have that

### v2

@@ Appendix A LCLs on graphs with sub-exponential growth
In this section, we show that on graphs of sub-exponential growth, we can solve any LCL problem in constant time by using $1$ bit per node of advice (we remind the reader that, as per definition of LCLs, the class of graphs that we consider are the ones with constant maximum degree, hence throughout this section we assume $\Delta=O(1)$). In more detail, we will prove the following theorem, and devote the rest of the section to proving it.
Theorem A.1.
@@
Let $\Pi$ be an LCL problem, and let $\mathcal{G}$ be a family of graphs of maximum degree $\Delta$ of sub-exponential growth in which $\Pi$ is solvable. Then, there exists a uniform fixed-length sparse $(\mathcal{G},\Pi,1,O(1))$-advice schema.
@@
We start by giving a formal definition of sub-exponential growth. For a node $v$, we denote with $N_{\leq d}(v)$ the set of nodes at distance at most $d$ from $v$, and we denote with $N_{=d}(v)$ the set of nodes at distance exactly $d$ from $v$.
Definition A.2 (Sub-exponential growth).
@@
A family $\mathcal{G}$ of graphs is of sub-exponential growth if, for any constant $c>0$, there exists a constant $x_{0}$ such that, for any $x\geq x_{0}$, for any $G=(V,E)\in\mathcal{G}$, for any $v\in V$, $|N_{\leq x}(v)|\leq 2^{c\cdot x}$.
@@
On a high level, we would like to exploit that, on graphs with sub-exponential growth, it is possible to somehow compute a clustering such that the number of nodes on the border of each cluster is much less than the number of nodes inside each cluster. Ideally, this property would allow us to encode the solution of the nodes of the border of a cluster by using the nodes inside that cluster. Then, inside each cluster, by brute force, we can complete the partial solution assigned to the border of the cluster, and hence each node would know its own part of the solution, as desired. However, things are not so simple: (1) not only we need to encode the solution of the nodes that are in the borders of the clusters, but we also need to be able to encode a clustering of the graph, and all this by using only one bit per node; (2) we need to make sure that the decoding process is able to decode the right thing, and this requires that the nodes that encode some information and that are inside two different clusters are far enough from each other. This latter issue forces us to prove some stronger properties of graphs with sub-exponential growth, that is that it is possible to compute a clustering such that the number of nodes that are very close to the center of the cluster is at least as large as the number of nodes in the border (Lemma˜A.3). In more detail, we prove that, for each node $v$, there exists some value $\alpha$ such that, for a cluster $C_{v}$ centered at $v$, the set $N_{\leq\alpha}$ of nodes of $C_{v}$ that are at distance at most $\alpha$ from $v$ is such that all nodes in $N_{\leq\alpha}$ are far enough from the border and they are at least as many as the ones in the border. We will use only nodes in $N_{\leq\alpha}$ to encode any useful information for us. In order to encode the clustering we do the following. First, we compute a distance-$d$ coloring of the graph (for some fixed $d$), that is, a coloring of the nodes such that nodes with the same color are at distance strictly larger than $d$. Then, we proceed through color classes in some order and at each phase $i$ we assign nodes to some cluster centered at a node with color $i$ (the nodes that will not be assigned to any cluster in this phase will be processed later). In this case, we say that $i$ is the color of the cluster. After we are done with color-class $i$, we remove the nodes that are part of some cluster and we repeat the procedure by considering the next color on the remaining subgraph. The distance coloring ensures that we avoid any kind of collisions with other centers of the same color. In order to encode the color of a cluster $C_{v}$ centered at $v$ we use a path that starts at node $v$ and is contained in $N_{\leq\alpha}$ (and hence it is far enough from the border of the cluster). Then, we use an independent set of the nodes in $N_{\leq\alpha}$ in order to encode the solution of the border (Lemma˜A.3 ensures that these nodes are enough for our purposes). The fact that we use only nodes in $N_{\leq\alpha}$ to encode information ensures that nodes that carry information regarding different clusters are far enough from each other (even if the clusters are adjacent) and hence we avoid wrong decodings. In order to be able to distinguish different encodings inside a cluster (e.g., the color of the cluster, the output of the border, the center of the cluster) we use the following scheme. Consider the connected components induced by the nodes having as input a $1$: the $1$s that form an independent set are the ones that encode the solution of the border; the other $1$s are used to encode the cluster center and its color, where four $1$s encode the center of the cluster, two $1$s encode a 0 and three $1$s encode a $1$.
@@
We start by proving a technical lemma regarding graphs of sub-exponential growth, which will be useful later.
Lemma A.3.
@@
Let $\mathcal{G}$ be a family of graphs of sub-exponential growth that have degree bounded by $\Delta$. For any integer $r>0$, let $c:=\log(1+1/\Delta^{r})/(3r)$, and let $x_{r}:=\max\{4r,x_{0}\}$, where $x_{0}$ is the value given by Definition˜A.2 for $\mathcal{G}$, when using parameter $c$. Then, for all $G=(V,E)\in\mathcal{G}$, and all $v\in V$, there exists a value $\alpha$ in $\{x_{r},\ldots,2x_{r}\}$ such that $|N_{\leq\alpha}(v)|\geq\Delta^{r}|N_{=\alpha+r}(v)|$.
Proof A.4.
@@
Assume, for a contradiction, that, for any $\alpha$ in $\{x_{r},\ldots,2x_{r}\}$, it holds that $|N_{\leq\alpha}(v)|<\Delta^{r}|N_{=\alpha+r}(v)|$. We prove, by induction, that, under this assumption, $|N_{\leq x_{r}+kr}(v)|\geq(1+1/\Delta^{r})^{k}\cdot|N_{\leq x_{r}}(v)|$, for any $k$ satisfying $0\leq kr\leq x_{r}$. For the base case $k=0$, we need to prove that $|N_{\leq x_{r}}(v)|\geq|N_{\leq x_{r}}(v)|$, which trivially holds. Assume that the statement holds for $k$, we show that it holds for $k+1$. We can lower bound $|N_{\leq x_{r}+(k+1)r}(v)|$ with $|N_{\leq x_{r}+kr}(v)|+|N_{=x_{r}+(k+1)r}(v)|\geq|N_{\leq x_{r}+kr}(v)|+|N_{\leq x_{r}+kr}(v)|/\Delta^{r}=|N_{\leq x_{r}+kr}(v)|(1+1/\Delta^{r})\geq(1+1/\Delta^{r})^{k}\cdot|N_{\leq x_{r}}(v)|\cdot(1+1/\Delta^{r})=(1+1/\Delta^{r})^{k+1}\cdot|N_{\leq x_{r}}(v)|$, proving the inductive step. Moreover, by the assumption of sub-exponential growth, we have that $|N_{\leq x_{r}+kr}(v)|\leq 2^{c(x_{r}+kr)}$. By combining the obtained bounds, we obtain the following:
$(1+1/\Delta^{r})^{k}\cdot|N_{\leq x_{r}}(v)|\leq|N_{\leq x_{r}+kr}(v)|\leq 2^{c(x_{r}+kr)}.$
@@
$(1+1/\Delta^{r})^{k}\cdot|N_{\leq x_{r}}(v)|>2^{c(x_{r}+kr)}.$
Since $|N_{\leq x_{r}}(v)|\geq 1$, it is sufficient to show that there exists a value of $k$ satisfying $0\leq kr\leq x_{r}$, such that
$(1+1/\Delta^{r})^{k}>2^{c(x_{r}+kr)},\text{ which is implied by }k>\frac{x_{r}}{\log(1+1/\Delta^{r})/c-r}.$
Since $c=\log(1+1/\Delta^{r})/(3r)$, we get that, in order to get our contradiction, it is sufficient to give an integer $k$ satisfying $k>x_{r}/(2r)$ and $0\leq kr\leq x_{r}$. For this purpose, we take $k:=\lceil\frac{x}{2r}\rceil+1$. Observe that $k>x_{r}/(2r)$ is clearly satisfied. Also, $kr\leq(\lceil\frac{x_{r}}{2r}\rceil+1)r\leq(\frac{x_{r}}{2r}+2)r=x_{r}/2+2r\leq x_{r}$, where the last inequality holds since $x_{r}\geq 4r$ (recall that $x_{r}$ is defined as $\max\{4r,x_{0}\}$).
@@
In order to show the desired advice schema, we first compute a clustering of the graph as follows. Let $r$ be a large-enough constant to be fixed later. Let $c$ and $x$ be the values given by Lemma˜A.3 as a function of $r$. Let $G\in\mathcal{G}$. At first, we compute a distance-$(5x)$ coloring on $G$. Since $G$ is a graph of sub-exponential growth, such a distance coloring can be done by using $2^{c\cdot 5x}$ colors (this is because $x\geq x_{0}$, where $x_{0}$ is the value given by Definition˜A.2 when using parameter $c$). Then, we process nodes by color classes, in ascending order, and, at each step $i$, we assign some nodes to some clusters. Let $G_{i}$ be the subgraph of $G$ induced by nodes that, at the end of step i1i-1, do not belong to any cluster yet. Let $v$ be a node of color $i$. We consider two possible cases: either, in $G_{i}$ it holds that $|N_{=2x}(v)|=0$, or in $G_{i}$ it holds that $|N_{=2x}(v)|>0$. For each node in the latter case, we compute a cluster as follows. Let $\alpha_{v}$ be the value of $\alpha$ given by Lemma˜A.3 for node $v$. We create a cluster containing $v$ and all the nodes of $G_{i}$, that, in $G_{i}$, have distance at most $\alpha_{v}+r$ from $v$. We call $v$ the center of this cluster. Observe that, at the end of this cluster-formation procedure, each node $v$ that does not belong to any cluster satisfies that, within distance $2x$, it sees the whole connected component containing $v$ induced by nodes that do not belong to any cluster.
@@
We provide an assignment of bits to the nodes of $G$ that encodes the clustering. We will later show how to add other bits on top of such an assignment and encode a solution of $\Pi$ in such a way that it is clear which bits encode the clustering and which bits encode the solution of $\Pi$. For each cluster $C$, let $v$ be its center, and let $i$ be the color of $v$. For any $u\in C$, let $d_{u}$ be the distance of $u$ from $v$. By the construction of $C$, there must exist $u\in C$ such that $d_{u}=y:=\lfloor x/2\rfloor$. Consider an arbitrary path $P_{v}=(v_{1},\ldots,v_{y})$ of length exactly $y$ where $v_{1}=v$ and $v_{y}=u$. This is the path that we will use for encoding the color of $v$. Let $B$ be the bit-string representation of $i$, which is of length $\lceil\log i\rceil$. We modify $B$ by replacing each 0 with the bit-string $110$, and each $1$ with the bit-string $1110$, obtaining the bit-string $B^{\prime}$. Then, we define $B^{\prime\prime}$ as the concatenation of $11110110$ with $B^{\prime}$ and with 0. In total, we use at most $4|B|+9=4\lceil\log i\rceil+9\leq 4\lceil\log(2^{c\cdot 5x})\rceil+9=20cx+9$ bits. By setting $r$ large enough, we obtain a small-enough $c$ and a large-enough $x$ such that $20cx+9\leq\lfloor x/2\rfloor=y=|P_{v}|$. Hence, we can assign the $j$th bit of $B^{\prime\prime}$ to the node $v_{j}$ of $P_{v}$. Finally, we assign the bit 0 to all nodes that did not get any bit from the previous bit-assignment. Note that all nodes within distance $\alpha_{v}+r$ (in $G_{i}$) from $v$ that are not part of $P_{v}$ have their bit set to 0.
@@
Proof A.6.
If $u$ and $w$ have the same color, then they are at distance at least $5x$, and hence all nodes in $N_{\leq\alpha_{u}}(u)$ are at distance at least $x$ from nodes in $N_{\leq\alpha_{w}}(w)$, and hence the property holds. Therefore, assume, w.l.o.g., that $u$ has color $i$ and $w$ has color $j$, for $i<j$. Then, in $G_{i}$, all nodes in $C_{w}\cap N_{\leq\alpha_{w}}(w)$ are at distance at least $r>0$ from all nodes in $C_{u}\cap N_{\leq\alpha_{u}}(u)$, since the cluster $C_{u}$ contains all nodes within distance $\alpha_{u}+r$ from $u$ in $G_{i}$. This implies that they are at distance strictly greater than zero also in $G$.
Corollary A.7.
@@
Proof A.8.
The proof follows from Lemma˜A.5 and by the fact that each path contains only nodes within distance $y<\alpha$ from the center of its cluster.
@@
Assume that we already decoded the clusters having centers of color $1,\ldots,i-1$. We now show how to decode clusters having centers of color $i$. Each node gathers its $2x$-radius neighborhood. In this way, each node $v$ knows which of the nodes within distance $2x$ from $v$ are part of some cluster with center of color at most i1i-1, and which nodes within distance $2x$ do not belong to any cluster yet. Hence, each node $v$ knows which nodes within distance $2x$ are part of $G_{i}$. In the following, all computation will be done with regard to $G_{i}$.
@@
  • for any $y+1\leq j\leq x$, nodes in $N_{=j}(v)$ have their bit set to 0;
  • there exists a path $P=(v_{1},\ldots,v_{y})$ satisfying that, for all $j$, the node $v_{j}$ is at distance exactly j1j-1 from $v$, and that contains all nodes in $N_{\leq y}(v)$ that have their bits set to $1$;
@@
  • nodes $v_{9},\ldots,v_{y}$} form a bit-string of the form $(110|1110)^{*}\,0\,0^{*}$.
By the encoding of the clusters, it is clear that all centers of color $i$ must satisfy these properties. Consider the set $S^{\prime}\subseteq S$ of nodes satisfying that, by decoding the bit-string on the path, they obtain color $i$. We prove that nodes in $S^{\prime}$ are all and only the centers of clusters of color $i$. Recall that, when we encoded clusters having a center $v$ of color $i$, we assigned a 0 to all nodes within distance $\alpha_{v}+r$ (in $G_{i}$) from $v$ that are not part of $P_{v}$. Hence, it is clear that, if a center $v$ has color $i$, then $v\in S^{\prime}$. Let $v$ be an arbitrary node in $S^{\prime}$, and let $P=(v_{1},\ldots,v_{y})$ be a path satisfying the properties above. Then we show that $v$ must be a center of color $i$. By Corollary˜A.7, nodes belonging to different paths form different connected components, and hence, if four nodes form a connected component containing four $1$s, then one of these nodes must be the center of a cluster. This implies that one node among $v_{1},\ldots,v_{4}$ must be the center of a cluster. By the above properties, we get that no node among $v_{2},v_{3},v_{4}$ can be in $S^{\prime}$, and hence $v_{1}=v$ is indeed a center of a cluster. Also, by the above properties, there are no other centers within distance $x$ from $v$. We therefore get that the bit-string encoded on $P$ is exactly the color of $v$, and hence that the color of $v$ is $i$.
@@
Let $\ell$ be a function mapping each node-edge pair of $G$ into a label of $\Pi$, such that the labeling obtained by $\ell$ is a solution for $\Pi$. We process clusters by colors in ascending order. Let $C_{v}$ be a cluster centered at $v$ with color $i$. In the following, all computation is done w.r.t. $G_{i}$. Let $\bar{r}$ be the checkability radius of $\Pi$. Let $S_{v}$ be the set of nodes within distance $\bar{r}$ (in $G_{i}$) from at least one node in $N_{=\alpha_{v}+r}(v)$. Note that $|S_{v}|\leq N_{=\alpha_{v}+r}(v)\Delta^{\bar{r}}$. Hence, by Lemma˜A.3, $|S_{v}|\leq|N_{\leq\alpha_{v}}(v)|\Delta^{\bar{r}}/\Delta^{r}$. Consider the labeling given by $\ell$ to all node-edge pairs incident to the nodes in $S_{v}$. We can encode such a labeling by using a bit-string of length $k|N_{\leq\alpha_{v}}(v)|\Delta^{\bar{r}+1}/\Delta^{r}$, where $k$ is a constant that depends on the alphabet of $\Pi$. Let $B$ be such a bit-string. Let $Y$ be the set containing the nodes of $N_{\leq\alpha_{v}}(v)$ that have their bit set to $1$, plus all their neighbors. Let $Z=N_{\leq\alpha_{v}}(v)\setminus Y$. The nodes that have their bits set to $1$ are at most $y=\lfloor x/2\rfloor$, and each such node has at most $\Delta-1$ neighbors that have a 0. Recall that $v$ has at least one node at distance exactly $2x$ (as otherwise $v$ would not be a cluster center), and that $N_{\leq\alpha_{v}}(v)$ contains all nodes within distance $\alpha_{v}\geq x$ from $v$. We get that there exists an injective function mapping nodes that have a $1$ to nodes in $Z$. Thus, $|Z|\geq|Y|/\Delta$, and hence $|N_{\leq\alpha_{v}}(v)|=|Z|+|Y|\geq|Y|/\Delta+|Y|=|Y|(1+1/\Delta)$, which implies $|Y|\leq|N_{\leq\alpha_{v}}(v)|/(1+1/\Delta)$. Thus, we get that $|Z|\geq|N_{\leq\alpha_{v}}(v)|-|Y|\geq|N_{\leq\alpha_{v}}(v)|/(\Delta+1)$. Hence, any independent set in $Z$ has size at least $|N_{\leq\alpha_{v}}(v)|/(\Delta+1)^{2}$. By picking $r$ large enough, we have that